# Survival Analysis with Adversarial Regularization

### v3

@@ Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent work shows that Neural Networks (NNs) can capture complex relationships in SA. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade model performance. To address this, we leverage NN verification advances to create algorithms for robust, fully-parametric survival models. We introduce a robust loss function and use CROWN-IBP regularization to handle computational challenges in the Min-Max problem. Evaluating our approach on SurvSet datasets, we find that our Survival Analysis with Adversarial Regularization (SAWAR) method consistently outperforms baselines under various perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI). This demonstrates that adversarial regularization enhances SA performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets up to 150% across all perturbation magnitudes.
@@ I Introduction
Survival Analysis (SA) models the time until an event occurs and has extensive applications. For example, SA has been used extensively in medicine for evaluating treatments in cancer patients and drug trials [1]. Early methods relied on linear parametric models [2], but recent advancements utilize Neural Networks (NNs) to handle complex time-to-event data [3].
A major challenge in Survival Analysis (SA) is developing models robust to dataset uncertainty and variability. This is particularly essential for clinical applications where treatment efficacy must be demonstrated across diverse patient demographics, institutions, and data acquisition settings [4, 5, 6]. Data uncertainty can arise from human errors, such as poor handwriting by physicians [7] and ineffective medical record-keeping [8], as well as inherent noise during data collection [9]. Many approaches assume perfect measurement of covariates, neglecting aleatoric uncertainty [10], which in SA is associated with noise in the log of measured time-to-event [11]. Further, adversaries can manipulate data to make models over- or under-confident [12]. Despite these issues and the sensitivity of NNs to input uncertainties, robustness techniques against data perturbations in SA remains underexplored.
To address these issues, this paper’s contributions include:
  • Survival Analysis with Adversarial Regularization (SAWAR) for training robust, fully-parametric, NN-based SA models using adversarial covariate perturbations and convex relaxations, enhancing model calibration and stability against noisy data.
  • Experiments demonstrating the sensitivity of existing NN-based SA methods to covariate perturbations across 10 publicly available medical SA datasets, highlighting an area for improvement in the current state of the art.
@@ II Related Work
This section reviews NN literature for SA, with more details in [13]. The first use of NNs in SA was a shallow Feedforward Neural Network (FFNN) to approximate the survival function [14]. Later, deeper models were developed, such as Cox-nnet [15], a NN for Cox Proportional Hazard (CPH) modeling [11]. Building on this, DeepSurv [16] introduced a CPH NN with dropout layers for better generalization in personalized treatment. Then, mixture models such as DeepWeiSurv [17] further advanced the field by using NNs to regress Weibull mixture distribution parameters. These methods enhance modeling capability but do not always improve model calibration and robustness.
Recent works focus on improving model calibration [18, 19, 20], which ensures predicted confidence matches true likelihood [21]. However, few works address adversarial robustness in SA, where small input perturbations should not mislead the model [22]. Although better calibration does not always enhance generalization [23], integrating adversarial robustness improves model trustworthiness in calibration and and stability on unexpected data [24].
Existing adversarial training methods for SA involve tradeoffs between model complexity, performance, and incorporating physical assumptions. Most techniques use Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples from a survival distribution non-parametrically [25, 26, 27]. These models implicitly define the survival distribution without assuming a functional/analytical form, offering increased flexibility to fit various data realizations [28, 29, 30]. However, this flexibility can violate Bayesian Occam’s Razor, which penalizes overly complex models [30]. To balance complexity and performance, inductive biases are used in explicit parametric models [29]. One common bias in SA is the memoryless property, satisfied by the exponential distribution [28]. Thus, this paper employs the exponential CPH model, widely used in SA [31, 32, 33], as well as in queuing theory [34], reliability analysis [3], and telecommunications [35].
@@ III Preliminaries on Survival Analysis
SA [11] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, the study happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring [36]. Such an SA dataset consists of $N$ independent instances/individuals, each denoted by index $i$, for which time between entry to a study and the subsequent event is recorded. Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that may affect the event of interest for the study) are measured. In summary, the dataset $\mathcal{D}$ has the following structure:
@@ III-A Cox Proportional Hazard Model
An exponential parametric survival model is expressed as
@@
$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta}$ (3)
@@
To incorporate the instance’s covariates, we use the prevalent CPH model [2]:
@@
where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative risk associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
The exponential CPH model characterizes the baseline hazard rate as $\lambda_{0}(t)=\lambda_{0}$. For simplicity, we assume an exponential baseline hazard and “absorb” the $\lambda$ term into the bias of the NN:
$\lambda_{0}e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (5)
The exponential CPH model’s PDF, CDF, and complement CDF are $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1.0-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively. $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
The population survival curve is the marginalization of the instance survival function over the covariates,
@@
Since the distribution of the covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate [18]:
@@
Given these definitions, the next section describes the baseline objective for finding model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
@@ III-B Baseline Objective Functions
We follow [38, 25] to define two objective functions based on negative log likelihood (Section III-B1) and the rank correlation (Section III-B2).
@@ III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the log of the time-to-event likelihood over $\mathcal{D}$ specified by the CPH model [11]:
@@
The first summation in Eq. 8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time of occurrence is observed. The second summation in Eq. 8 is the log likelihood when right censoring occurs.
@@ III-B2 Ranking Objective
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [38, 39]. We set $\sigma=1$ for all experiments.
@@ III-B3 Combined Objective
The combined right censored negative log likelihood objective and the risk ranking objective, with risk ranking objective weight trade-off $w=\frac{1}{\text{batch\_size}}$, is
$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}).$ (10)
This combined objective provides a trade-off between finding the NN parameters $\boldsymbol{\theta}$ that maximize the likelihood of observing $\mathcal{D}$ and maximize a rank correlation between the instance risk scores and the observed failure times. We will denote the combined loss as the baseline objective. Our baseline model’s objective is similar to Deep Regularized Accelerated Failure Time (DRAFT) [25] but with a different ranking objective function. Thus, we consider the baseline training method for robust SA baseline method as DRAFT.
Thus far, we have described an objective function that improves predictive performance:
@@
In the following sections, we address how to develop NN models that are robust against adversarial perturbations.
@@ IV Preliminaries on Adversarial Training
In practice, the Min-Max problem posed in [42] is intractable to compute exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 12, leading to models with varying degrees of robustness to different data uncertainties. To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
IV-A Certifiable Robustness
One such approximation is to use techniques from NN verification during training. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data [40, 43]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [44, 40] showed the linear relaxation can be extended to the entire objective function:
@@ IV-A Certifiable Robustness
This relaxation is convex with respect to NN weights, enabling efficient optimization, and by optimizing over an upper bound on the inner maximization, it accounts for a strong adversary during training.
@@ IV-B Adversarial Perturbations
Another way to approximate Eq. 20 is to use adversarial perturbations. Whereas CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound:

### v4

@@ Abstract
Survival Analysis (SA) models the time until an event occurs, with applications in fields like medicine, defense, finance, and aerospace. Recent research indicates that Neural Networks (NNs) can effectively capture complex data patterns in SA, whereas simple generalized linear models often fall short in this regard. However, dataset uncertainties (e.g., noisy measurements, human error) can degrade NN model performance. To address this, we leverage advances in NN verification to develop training objectives for robust, fully-parametric SA models. Specifically, we propose an adversarially robust loss function based on a Min-Max optimization problem. We employ CROWN-Interval Bound Propagation (CROWN-IBP) to tackle the computational challenges inherent in solving this Min-Max problem. Evaluated over 10 SurvSet datasets, our method, Survival Analysis with Adversarial Regularization (SAWAR), consistently outperforms baseline adversarial training methods and state-of-the-art (SOTA) deep SA models across various covariate perturbations with respect to Negative Log Likelihood (NegLL), Integrated Brier Score (IBS), and Concordance Index (CI) metrics. Thus, we demonstrate that adversarial robustness enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization across diverse datasets by up to 150% compared to baselines.
@@ I Introduction
Survival Analysis (SA) models the time until an event occurs with extensive applications in various fields including medicine, system reliability, economics, and marketing. Medical research uses Survival Analysis (SA) to evaluate treatment efficacy and illness progression [1]. Engineers assess a system’s reliability with SA to improve design and maintenance [2]. Economists and social scientists use SA (also known as duration analysis) to investigate the length of unemployment spells and other major life events [3]. Marketing uses SA to aid in the examination of customer retention and churn rates [4]. While many of these domains have traditionally relied on linear parametric SA models [5], recent advancements have leveraged Neural Networks (NNs) to handle such complex time-to-event data [6]
A major challenge in SA is developing models that effectively handle complex time-to-event data while also ensuring robustness against dataset uncertainty, accurate calibration, and strong generalization. For instance, data uncertainty in medical applications can arise from various sources, including human factors such as illegible handwriting by physicians [7] and poor medical record-keeping [8], as well as inherent noise from medical equipment during data collection [9], termed as aleatoric uncertainty. Most of the existing approaches address aleatoric uncertainty by assuming noise in the logarithm of measured time-to-event data [10]. However, they often overlook uncertainty originating from the model’s input covariates.
SA models ought to generalize over various input covariates to sustain high accuracy in real-life applications. In medical studies, this entails high accuracy over diverse patient demographics, institutions, and data acquisition settings. While traditional SA models, such as Linear Cox-Weibull model [5], offer inductive biases specific to time-to-event analysis that aid generalization, they struggle with complex data relationships. On the other hand, while NN models can manage complex data, they may not generalize well to unseen data [11]. Combining the inductive biases of traditional SA models with NN techniques—such as modeling intensity or hazard functions with NNs —can both improve generalization and handle complex data [12]. Nonetheless, NN in SA may still suffer from poor model calibration, even with good generalization.
NNs are particularly sensitive to input perturbations [13]. Adversaries can exploit this sensitivity to manipulate data, causing models to make over- or under-confident predictions [14, 15]. Additionally, data perturbations may be intentionally introduced with techniques like differential privacy to protect sensitive information, such as personally identifiable information of patients [16]. Although better calibration does not always enhance generalization [17], integrating adversarial robustness to input covariate perturbation can jointly improve model stability, trustworthiness, and calibration when encountering unseen data [18].
Despite these challenges and the sensitivity of NNs to input uncertainties, robustness techniques for handling data perturbations in NN-based SA remain underexplored. To address these challenges, we make the following contributions:
  • We propose a novel training objective, called Survival Analysis with Adversarial Regularization (SAWAR), designed to create robust, fully-parametric NN-based SA models by incorporating adversarial covariate perturbations via convex relaxations, and thereby improve model calibration and generalization against noisy data.
  • Extensive experiments across 10 publicly-available SA datasets reveal the sensitivity of existing NN-based SA models to covariate perturbations with respect to Concordance Index (CI), Integrated Brier Score (IBS), and Negative Log Likelihood (NegLL) metrics, as well as survival curves.
  • SAWAR enhances SA predictive performance and calibration, mitigating data uncertainty and improving generalization by up to 150% compared to baseline adversarial training methods and state-of-the-art (SOTA) NN-based SA models.
@@ II Related Work
The integration of NNs into SA began with initial models that replaced traditional linear approaches. Early works used a shallow Feedforward Neural Network (FFNN) to replace the linear estimate of the log-hazard in a Cox Proportional Hazard (CPH) model [19]. Building on this, Cox-nnet [20] substituted the input covariates of the linear log-hazard function of a CPH model [21] with a deeper NN. Subsequently, DeepSurv [22] further advanced this approach by replacing the entire linear estimate with a deep NN and adding dropout to improve generalization. The trend continued with mixture models like DeepWeiSurv [23], which used NNs to regress Weibull mixture distribution parameters, offering increased modeling flexibility while maintaining traditional statistical model biases.
As NN models evolved, the SA field saw the introduction of non-parametric approaches. For instance, DeepHit utilized NNs to learn the distribution of survival times directly without assumptions about the underlying stochastic process [24]. Even more flexible techniques emerged, such as those employing Conditional Generative Adversarial Networks (CGANs) to generate time-to-event samples non-parametrically [25, 26, 27], with CGAN serving as an implicit likelihood [28]. This approach defines the survival distribution without assuming a specific functional form [29, 30]. For a more detailed overview of deep learning methods for SA, we refer the readers to [12].
Despite these advances, integrating NNs with traditional SA models has been limited in addressing model calibration and robustness as discussed in Section I. Recent research has thus focused on improving model calibration [31, 32, 33] to ensure that predicted confidence levels reflect true likelihoods [34]. However, better calibration does not necessarily enhance generalization [17]. On the other hand, integrating adversarial robustness has been shown to improve both calibration and generalization when dealing with unexpected data [18]. Nonetheless, there remains a gap in addressing adversarial robustness in SA to ensure that where small input perturbations do not mislead model predictions [15].
Existing adversarial training methods for SA often face trade-offs between model complexity, prediction performance, and the incorporation of physical assumptions. While Generative Adversarial Network (GAN)-inspired methods incorporate adversarial components through a minimax game between the generator and discriminator [35, 36], they typically exhibit training instability and may violate Bayesian Occam’s Razor that advocates simpler models [30]. To balance complexity, interpretability, and generalization, several methods adopt time-to-event inductive biases through specific functional forms of SA models with deep NN and/or non-parametric approaches [22, 20, 23, 37].
One such inductive bias in SA is the memoryless property, satisfied by the exponential distribution [29]. This paper employs the exponential CPH model (c.f. Section III), as it is widely used in SA [38, 39, 40], as well as in related fields like queuing theory [41], reliability analysis [6], and telecommunications [42]. Crucially, this fully-parametric approach allows us to create an adversarially robust NN-based SA model that outperforms SOTA NN-based SA models including such as Deep Regularized Accelerated Failure Time (DRAFT) [25] and Adversarial Autoencoder with a Deep Cox Proportional Hazard (AAE-Cox) [37]. AAE-Cox aims to balance modeling flexibility and inductive biases by combining adversarial GAN autoencoders with Deep Cox proportional networks. AAE-Cox has been shown to outperform methods like DeepSurv, TCAP [43], and DCAP [44].
@@ III Preliminaries on Survival Analysis
Figure 1: Visual of time-to-event dataset with exact event times and right censoring. The solid circle on each patient line indicates when the event occurred for a given patient.
SA [21] is a statistical approach used to analyze and model the time $t\in\mathcal{R_{++}}$ until an event $e\in\{0,1\}$ happens (e.g., a disease occurs, a device breaks). Typically, a study of such events happens over a finite time period and instances/individuals may leave the study early before an event is recorded, with no follow-up. In those cases, the dataset records that no event occurred up until the exit time of the study, which is called right censoring  [45]. Such an SA dataset consists of $N$ independent instances, each denoted by index $i\in[1,\ldots,N]$, for which the time between entry to a study and the subsequent event is recorded. Upon entry time to a study, each instance’s covariates $\boldsymbol{x}_{i}\in\mathcal{R}^{d}$ (i.e., variables that govern the event of interest) are measured. A visual of time-to-event data with right censoring is shown in Fig. 1. In summary, the dataset $\mathcal{D}$ has the following structure:
@@ III-A Cox Proportional Hazard Model
An exponential SA model has the parametric form:
@@
$\lambda(t)=\frac{f(t)}{S(t)}=\lim_{\Delta\rightarrow 0}\frac{P[t<T<t+\Delta]}{\Delta},$ (3)
@@
To incorporate each instance’s covariates in predicting failure rate, we use the prevalent CPH model [5]:
@@
where $\lambda_{0}(t)$ is a baseline hazard, $e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}$ is the relative hazard associated with covariates $\boldsymbol{x}$, and $G_{\boldsymbol{\theta}}(\boldsymbol{x})$ is an NN with parameters $\boldsymbol{\theta}$ in our setting.
The exponential CPH model assumes a time-independent baseline hazard rate $\lambda_{0}(t)=\lambda_{0}$. This allows us to “absorb” the $\lambda_{0}$ term into the bias of the NN:
$\lambda_{\boldsymbol{\theta}}(t,\boldsymbol{x})=e^{\log\lambda_{0}+G_{\boldsymbol{\theta}}(\boldsymbol{x})}=e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}.$ (5)
Further, the exponential CPH model’s PDF, CDF, and complement CDF are given by $f_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{G_{\theta}(\boldsymbol{x})}e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, $F_{\boldsymbol{\theta}}(t|\boldsymbol{x})=1-e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$ , $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})=e^{-e^{G_{\boldsymbol{\theta}}(\boldsymbol{x})}t}$, respectively. $S_{\boldsymbol{\theta}}(t|\boldsymbol{x})$ is also known as the instance survival function.
The population survival curve is defined as the marginalization of the instance survival function over the covariates:
@@
Since the distribution of covariates $p(\boldsymbol{x})$ is typically unknown, the integral in Eq. 6 is often replaced with a Monte Carlo estimate over a dataset $\mathcal{D}$ [31]:
@@
Given these definitions, the next section describes the baseline objective for estimating model parameters $\boldsymbol{\theta}$ from dataset $\mathcal{D}$.
@@ III-B Baseline Objective Functions
We employ an objective function that combines negative log likelihood (Section III-B1) and rank correlation (Section III-B2), following [47, 25].
@@ III-B1 Right Censored Log Likelihood Objective
The $\mathcal{L}_{LL}$ objective is the logarithm of the time-to-event likelihood over $\mathcal{D}$ as specified by the CPH model [21]:
@@
The first summation in Eq. 8 is the log likelihood of the datapoints corresponding to when the event occurs and the exact time that occurrence is observed. The second summation in Eq. 8 is the log likelihood when right censoring occurs, where it is known that no event occurred at least up until the censoring time.
@@ III-B2 Ranking Objective
where $\eta(\mathbf{x},\mathbf{y})=\exp(-\frac{(\mathbf{x}-\mathbf{y})}{\sigma})$ and $A_{i,j}$ indicates an “acceptable” comparison pair [47, 48]. Two instances are comparable if the instance with lower observed time experienced an event, i.e., if $A_{i,j}=\mathds{1}(t_{i}<t_{j}\cap e_{i}=1)$. We set $\sigma=1$ for all experiments. This objective is a continuous differentiable proxy for the non-continuous non-differentiable concordance index [49].
@@ III-B3 Combined Objective
The combined objective over the right censored dataset $\mathcal{D}$ is:
$\mathcal{L(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})}=-\mathcal{L}_{LL}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e})+w\cdot\mathcal{L}_{rank}(\boldsymbol{\theta};\boldsymbol{X},\boldsymbol{t},\boldsymbol{e}),$ (10)
where $w$ is the hazard versus ranking objective weight trade-off. Accordingly, $\boldsymbol{\theta}$ is estimated by solving the following optimization problem:
@@
This combined objective provides a trade-off between maximizing the likelihood of observing $\mathcal{D}$ and maximizing a rank correlation between the instance hazard scores and the observed failure times. From here on, we term this combined loss as the baseline objective. We note that an NN-based SA model DRAFT [25] also employs a combined objective. Our baseline objective incorporates a different ranking objective than DRAFT and outperforms DRAFT as discussed in Section VI.
@@ IV Preliminaries on Adversarial Training
In practice, the Min-Max problem posed in [13] is intractable to solve exactly. In particular, the inner maximization requires finding a global optimum over a high dimensional and non-convex loss function over input covariates with the current NN parameters. Therefore, many recent adversarial training methods propose approximations to Eq. 12, leading to models with varying degrees of robustness to different data uncertainties. To understand the robustness properties of SA models to a variety of data uncertainties present in real applications, this paper investigates several perturbation methods, described next.
IV-A Verifiable Robustness
One such approximation is to use techniques from NN verification during training. Neural network robustness verification algorithms determine the output neuron’s upper and lower bounds for all inputs in a set $\mathcal{B}$, typically constrained by a norm-bounded perturbation [51]. For example, CROWN-Interval Bound Propagation (CROWN-IBP) combines CROWN and Interval Bound Propogation (IBP) methods to obtain a convex relaxation for the lower bound and upper bound of each layer in the NN with respect to input data within the set $\mathcal{B}$. [50, 51, 52]. CROWN-IBP is commonly used for verifying NN properties over a range of possible inputs. [53, 50] showed the linear relaxation can be extended to the entire objective function:
@@ IV-A Verifiable Robustness
[…]
@@ IV-B Adversarial Perturbations
Another way to approximate Eq. 12 is to use adversarial perturbations.
IV-B1 Fast Gradient Sign Method (FGSM)
While CROWN-IBP provides an upper bound on the inner maximization (i.e., $\max_{\boldsymbol{\tilde{X}}}\mathcal{L}(\theta;\boldsymbol{\tilde{X}})$), adversarial perturbation algorithms typically provide a lower bound: